# The Theory of Economic Complexity

### v1

@@ Abstract
Economic complexity estimates rely on eigenvectors derived from matrices of specialization to explain differences in economic growth, inequality, and sustainability. Yet, despite their widespread use, we still lack a principled theory that can deduce these eigenvectors from first principles and place them in the context of a mechanistic model. Here, we calculate these eigenvectors analytically for a model where the output of an economy in an activity increases with the probability the economy is endowed with the factors required by the activity. We show that the eigenvector known as the Economic Complexity Index or $ECI$ is a monotonic function of the probability that an economy is endowed with a factor, and that in a multi-factor model, it is an estimate of the average endowment across all factors. We then generalize this result to other production functions and to a short-run equilibrium framework with prices, wages, and consumption. We find that our main result does not depend on the introduction of prices or wages, and that the derived wage function is consistent with the convergence of economies with a similar level of complexity. Finally, we use this model to explain the shape of networks of related activities, such as the product space and the research space. These findings solve long standing theoretical puzzles in the economic complexity literature and validate the idea that metrics of economic complexity are estimates of an economy being endowed with multiple factors.
@@ 1 Introduction
For about two decades, scholars working in economic complexity have focused on two key goals. The first one, is the development of economic complexity metrics to explain international and regional variations in economic growth [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], inequality [20, 21, 22, 23, 24, 25, 26, 27, 28], and sustainability [29, 30, 31, 32, 33, 34, 35, 36, 37, 38].111among other outcomes [39, 40, 41, 42, 43, 44, 45, 46] The second one, is mapping networks of related activities, such as the product space [47, 2], industry space [48, 49, 50], or research space [51, 52], to explain path dependencies in economic development[53]. Yet, despite several attempts to develop a mathematical theory of economic complexity [1, 2, 54, 13, 12, 55, 56, 57, 58, 59, 60, 61, 62, 63], we still lack an analytical connection between the metrics used in the empirical literature and a production function based model that can provide a clear interpretation of these metrics in terms of a model’s assumptions.222for a review of the field see [64, 65]
Here, we connect empirical economic complexity work with a few theoretical models to provide four contributions.
@@
First, we calculate the eigenvector known as the economic complexity index ($ECI$) for a capability model that expands Kremer’s O-Ring model model of development [66]. In our model, economies (such as countries or cities) are endowed with a probability of having the capabilities required by each activity (such as products or industries). This means the the output of economies is constrained by the capabilities they have and the geography of activities is limited to the places endowed with the capabilities they require. We solve the one capability instance of this model analytically to show that the economic complexity index, or $ECI$, divides economies among those with an above and below average probability of having the capability. Interestingly, this property is independent of how capabilities are distributed and can be generalized to other production functions (such as a shifted Cobb-Douglas factor intensity function).
@@
Second, we extend this result numerically to models involving many capabilities assigned idiosyncratically to each economy. We show that in this case $ECI$ is a monotonic function of the average capability endowment of an economy. By exploring models combining correlated and uncorrelated capabilities, we show this result to be robust to substantial levels of noise, holding even when more than 50 percent of an economy’s capability endowments are random. This helps show that $ECI$ is indeed a measure of economic complexity, since it is a statistic informing us if an economy is endowed with multiple capabilities.
@@
Third, we extend the single capability model to a short-run equilibrium model and calculate wages, prices, and consumption. We show analytically that under these assumptions $ECI$ still separates economies among those with high and low capability endowments. We also determine an equilibrium wage to help interpret the known empirical relationship between economic complexity and growth, and show that the prices of goods in this model follows a concave function of their capability requirements.
@@
There are a few reasons why these results should interest those working on economic complexity and international development.
@@
First, while the economic complexity index or $ECI$ enjoys wide adoption in policy circles [^3] For example, it is the number one mission of Malaysia’s New Industrial Master Plan [68], it was used in the recent competitiveness report by Mario Draghi [69], and it is a key development target for rich resource intensive economies, such as Saudi Arabia and the United Arab Emirates. It has also motivated the creation of regional reports for Australia[70], Turkey [71], Uruguay [72], Russia[73, 74], Mexico [75, 76], Quebec[77], and Italy[78], among other places., the lack of a theoretical foundation has left them open to critics of being an ad-hoc or uninterpretable measure [79, 13, 80, 61]. Our findings provide a clear interpretation of $ECI$ in the context of multi-factor model of production. We show that $ECI$ provides a monotonic estimate of the factor endowment of an economy derived from a multi-product specialization matrix. This provides an interpretation in terms of a model’s parameters that is consistent with previous work exploring the interpretability of the economic complexity index as a clustering method [56, 60, 59] and the connection between $ECI$ and log-supermodularity [55].
@@
Second, these findings dispel the notion that economic complexity is a measure of diversity, as it was originally suggested [1]. The analytical solutions show that economies specialized in the largest number of activities (the more diverse economies) are not necessarily the ones with the highest probability of having a capability [^4] The notion that economic complexity is different from diversity was noted theoretically by [56] and has been in the literature from the beginning, since the work introducing $ECI$ showed that measures of diversity or concentration, such as entropy or the Herfindahl-Hirschmann index (HHI) failed to explain future economic growth as $ECI$ did [1].. In fact, the model predicts that economic development is a process of diversification only until a certain point, since economies with the highest capability endowments specialize in complex activities–and are therefore–less diverse than slightly less complex economies. This also provides a theoretical foundation for the finding that countries at high-level of development tend to specialize (e.g., Imbs and Warcziag [81]), which is consistent with the notion that $ECI$ is higher for “small” yet sophisticated economies, such as those of Singapore, Switzerland, and Finland, while larger and more diverse economies, like those of Spain and Italy, are not necessarily as complex. Still, the model predicts a positive correlation between capability endowments and diversity, but through a non-monotonic function, explaining why measures of diversity or concentration are non-ideal estimates of the complexity of an economy.
@@
Third, these results also provide a mean to interpret the structure of the networks of related activities, such as the product space [47], industry space [48], or research space [51]. These networks have been used extensively to model path dependencies and generate measures of export or employment potential [53, 82, 51, 48, 50, 49, 83, 64, 84, 85, 86, 87, 88, 89]. Yet, the structure of these networks differs depending on the data used to generate them. For instance, product space networks derived from exports data, are known to have a core composed of densely interconnected activities with are high in complexity [47]. Research spaces, connecting academic fields based on citations [67] or co-authorships [51], follow a ring structure, with fields connected with a few neighbors and without a clear center [^5] In simple, the ring: medicine, biology, chemistry, physics, computer science and math, economics, cognitive science, neuroscience, and back to medicine. While these differences in structure are self-evident, we hitherto lack a way to explain them based on the mechanics of a model. Here, we show how to generate these structures by changing the shape of the capability endowment matrices.
@@
Together, these findings help solve some long-standing puzzles in the economic complexity and international development literature by providing a bridge linking its theoretical and empirical contributions.
@@ 1.1 Empirical and Theoretical Work in Economic Complexity
Empirical work in economic complexity usually starts with matrices summarizing the geography of many economic activities (e.g., exports by country and product, payroll by city and industry, patents by city and technology, etc.). These rectangular matrices (or bipartite networks) are then used for two things. The first one is to estimate networks of similar activities [90, 82, 51, 47, 53, 64, 48, 91, 92] which are used to create measures of diversification potential or relatedness. These measures are based on the notion that economies are more likely to enter (and less likely to exit) activities that share capabilities with each other.
@@
The second one is to create measures of an economy’s structure, or of the value of the portfolio of activities an economy specializes in, known as measures of economic complexity [1, 64, 11, 93, 2, 12, 13, 94]. These measures were also motivated as estimates of the capabilities available in an economy [1] and are often based on the assumption that high-complexity economies specialize in high-complexity activities. In fact, the definition of economic complexity introduced by [1], the economic complexity index of $ECI$, defines the complexity of an economy as the average complexity of the activities it specializes in, and the complexity of an activity as the average complexity of the economies specialized in that activity. [^6] A similar definition was proposed over a decade later by [13]. In their words: “If a country is known to be more capable than another, say the United States (US) versus Bangladesh (BG), then one can identify any good $k$ as more complex than another reference good $k_{0}$ if, relative to the reference good, it is more likely to be exported by the United States than Bangladesh. [$\dots$] Conversely, if a good is known to be more complex than another, say medicines (ME) versus men’s underwear (UW), then one can identify any country $i_{1}$ as more capable than another reference country $i_{0}$ if, relative to the reference country, it is more likely to export medicines than underwear. ”. These measures of complexity, in particular $ECI$ enjoy wide adoption in international and regional development circles, as they have been shown to be robust estimators of future economic growth [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], and of international variations in inequality [20, 21, 22, 23, 24, 28, 35], and emissions [29, 30, 31, 32, 33, 34, 95, 36, 37]. This supports the notion that an economy’s pattern of specialization matters for subsequent economic development [96, 97], which was a key policy intuition motivating these quantitative efforts [^7] This policy intuition is connected to an old debate in development economics, going back to at least Alexander Hamilton’s Report on Manufactures [98], which advocated for the industrialization of the United States, and has been central to the works of scholars such as Rosenstein-Rodan [99, 100], Rostow[101], Hirschman[102], Prebisch[103], Gerschenkron[104], and Balassa[105]. For a discussion on how these different development theories related to economic complexity see [106]. Yet, despite copious empirical work, we still lack an understanding of why the eigenvectors of these matrices are a good metric of an economy’s complexity and a good predictor of its subsequent development.
@@
Theoretical work on economic complexity has focused instead on the construction of models of economic development and innovation that follow a combinatorial tradition  [107, 108, 1, 66, 54, 109, 110, 111, 112, 113, 114, 115]. This tradition builds on the notion that economies are endowed with capabilities [116, 117, 118, 119, 2, 120], or factors, which activities may or may not require. Since these capabilities are complementary, the producing an activity requires the simultaneous presence of many of them. That’s why these theories have been dubbed as the “Lego” or “Scrabble” theory of development. In these models, the ability of economies to produce a product depend on having the right combination of capabilities, like in a proverbial game of scrabble where products are “words” and economies are endowed with “letters.”
@@
Here we build on the combinatorial model introduced by [1], which is a generalization of Kremer’s O-Ring model of development [66].
@@
The O-Ring model assumes a multi-step production process where the output of an economy is the product of the probabilities that it succeeds at each step. In other words, producing an item in the O-Ring model requires a sequence of tasks, each of which has a probability of failing. This implies that the output of an economy decays exponentially with the length of the production chain at a rate determined by the probability of succeeding at a task. The key outcome is that economies with higher probabilities of completing a task should specialize in activities requiring multiple steps888see also [121] for an extension of the O-Ring model to trade.. Here we focus on a generalized version of this model, where economies are endowed with probabilities of having a capability (similar to the probability of succeeding at a task in the O-Ring model) and where activities also differ in the probability of requiring a capability. This allows us to model matrices involving an arbitrary number of economies, activities, and capabilities, while also making the capabilities specific to activities and economies. The resulting matrices, which can be made equivalent in size to the ones used in the empirical literature, can be used to create theoretical estimates of the economic complexity eigenvector $ECI$ that we can connect to the key parameter in the model: the probability that an economy is endowed with a capability.
@@
We find that, for a wide variety of model specifications, $ECI$ recovers the key parameter in the model: the probability that an economy is endowed with a capability. This result is valid for models involving multiple capabilities, even when these are highly idiosyncratic, since $ECI$ recovers the average of the probability that an economy is endowed with a capability. We then generalize this finding to a Cobb-Douglas type factor intensity production function and use this to show that the ability of $ECI$ to separate among better and worse endowed economies can be generalized to any shifted production function of the form $Y_{cp}=B+f_{c}g_{p}$ where $f_{c}$ is a general function characterizing an economy and $g_{p}$ is a general function characterizing an activity.
@@
As in Kremer’s O-Ring model [66], our initial exercise builds on a supply-side model that assumes prices are exogenous and does not provide an explicit model of wages or demand. So, we then embed our model in a short-run equilibrium framework and estimate functions for the implied wages, consumption, and prices. We show that wages increase with the capability endowment of an economy, consumption grows with income, and prices are higher for more demanding products (products having a higher probability of requiring a capability). Surprisingly, our main result (that the economic complexity eigenvector separates among high- and low-capability economies) is unaffected by the introduction of these additional assumptions.
@@
Finally, we use this model to explore the connection between the structure of networks of related activities, such as the product space and research space, and show that it is possible to generate networks with a similar structure than the ones observed in the empirical literature by manipulation the capability endowment and requirements of economies and activities.
@@
The remainder of the paper is organized as follows. The next section (Section 2) introduces the single-capability model and solves it analytically. Section 3 generalizes these results numerically to several versions of a multi-capability model. Section 4 explores additional production functions and Section 5 embeds the model in a short-run equilibrium framework. Section 6 uses the model to explain the structure of networks of related activities, and the Section 7 concludes.

### v5

@@ Abstract
Economic complexity methods have become popular tools in economic development, economic geography, and innovation. Yet, despite their widespread adoption, we lack a mechanistic model that provides these methods with a solid mathematical foundation. Here, we analytically derive the economic complexity eigenvector associated with a mechanistic model where the output of an economy in an activity (e.g. of a country in a product) depends on the combined presence of the factors required by the activity. Using analytical and numerical derivations we show that the economic complexity index (or $ECI$) is a monotonic function of the probability that an economy is endowed with many factors, validating the idea that it is an agnostic estimate of the presence of multiple factors of production. We then generalize this result to other production functions and to a short-run equilibrium framework with prices, wages, and consumption, showing that the derived wage function is consistent with economies converging to an income that is compatible with their complexity. Finally, we use this model to explain differences in the networks of related activities, such as the product space and the research space, showing that the shape of these networks can be explained by different factor distributions. These findings solve long standing puzzles in the economic complexity literature and validate commonly used metrics of economic complexity as estimates of the combined presence of multiple factors.
@@ 1 Introduction
A key tenet of the economic complexity literature is the idea that the combined presence of factors of production can be estimated without having to define them. This notion is central to the two key contributions that jump-started the study of economic complexity during the beginning of this century.
The first example is the product space [1], a network of related products based on the idea that “if two goods are related because they require similar institutions, infrastructure, physical factors, technology, or some combination thereof, they will tend to be produced in tandem.” By using an outcomes based measure, the product space can be used to create estimates of economic potential that do not rely on defining specific factors of production, but that leverage instead implicit information about unknown factors present in patterns of co-specialization. Networks of related activities, such as the product space [1, 2], industry space [3, 4, 5], research space [6, 7], and technology space [8, 9], have become important tools in economic geography, innovation studies, and international development, as they provide a means to formalize notions of path dependency and estimate the chances that an economy is endowed with the factors needed to engage in an activity.
@@
The second example is the development of metrics of economic complexity, which attempt to estimate the combined presence of factors available in an economy. In [10], economic complexity metrics were introduced as a mean to estimate capabilities that are not directly observed or named. These metrics of complexity have also become useful tools in economic geography, international development, and innovation, because of their ability to explain international and regional variations in economic growth [10, 2, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], inequality [29, 30, 31, 32, 33, 34, 35, 36, 37], and sustainability outcomes [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. [^1] Among other outcomes [48, 49, 50, 51, 52, 53, 54, 55].. Yet, despite several attempts to develop a mathematical theory of economic complexity [10, 2, 56, 21, 20, 57, 58, 59, 60, 61, 62, 63, 64, 65], we still lack an analytical connection between the metrics used in the empirical literature and a production function based model from where we derive these metrics from first principles.  [^2] For a review of the field see [66, 67]. Here, we connect the empirical economic complexity literature with a theoretical model to provide four contributions.
@@
First, we provide an analytical derivation of the eigenvector known as the economic complexity index ($ECI$) for a model where economies (such as countries or cities) are differently endowed with the capabilities required by each activity (such as products or industries). In this model, the output of an economy is constrained by the capabilities it has while the geography of an activity is limited to the places endowed with the capabilities the activity requires. We solve the one capability instance of this model analytically and show that the economic complexity index, or $ECI$, is a vector separating economies with an above average capability endowment from those with a below average capability endowment. Interestingly, this property is independent of how capabilities are distributed and can be generalized to other production functions (such as a shifted Cobb-Douglas factor intensity function).
@@
Second, we extend this result numerically to models involving many capabilities assigned idiosyncratically to each economy. We show that in this case $ECI$ is a monotonic function of the average capability endowment of an economy and recovers the first singular vector of the matrix of capability endowments. By exploring models combining correlated and uncorrelated capabilities, we show this result to be robust to substantial levels of noise, holding even when more than 50 percent of the capability endowments are assigned at random. This helps show that $ECI$ is a measure of economic complexity, as it can captures whether an economy is endowed with multiple capabilities without having to make assumptions about their nature.
@@
Third, we extend the single capability model to a short-run equilibrium framework where we calculate wages, prices, and consumption. We show analytically that under these assumptions $ECI$ still separates economies among those with high and low capability endowments. We also determine an equilibrium wage to help interpret the known empirical relationship between economic complexity and growth, and show that the prices of goods in this model follows a concave function of their capability requirements, indicating a high premium for the production of complex goods.
@@
There are a few reasons why these results should be of interest to those working on economic complexity, economic growth, and international development.
@@
First, while the economic complexity index or $ECI$ enjoys wide adoption in policy circles [^3] For example, it is the number one mission of Malaysia’s New Industrial Master Plan [69], it was used in the recent European competitiveness report by Mario Draghi [70], and it is a key development target for rich resource intensive economies, such as Saudi Arabia and the United Arab Emirates. It has also motivated the creation of regional reports for Australia [71], Turkey [72], Uruguay [73], Russia [74, 75], Mexico [76, 77], Quebec [78], and Italy [79], among other places., the lack of a theoretical foundation has left it open to criticism of being an ad-hoc or uninterpretable measure [80, 21, 81, 63]. Our findings provide a clear interpretation for $ECI$ in the context of multi-factor model of production. We show that $ECI$ provides a monotonic estimate of the factor endowment of an economy derived from a multi-product specialization matrix. This provides an interpretation in terms of a model’s parameters that is consistent with previous work exploring the interpretability of the $ECI$ as a clustering method [58, 62, 61] and connecting $ECI$ with the notion of log-supermodularity [57].
@@
Second, these findings dispel the notion that economic complexity is a measure of diversity, as it was originally suggested [10]. The analytical solutions show that economies specialized in the largest number of activities (the more diverse economies) are not necessarily the ones with the highest probability of having a capability [^4] The notion that economic complexity is different from diversity was noted theoretically by [58] and has been in the literature from early on, since the work introducing $ECI$ showed that measures of diversity or concentration, such as entropy or the Herfindahl-Hirschmann index (HHI) failed to explain future economic growth as $ECI$ did [10].. In fact, the model predicts that economic development is a process of diversification only until a certain point, since economies with the highest capability endowments are expected to specialize in complex activities–and are therefore–less diverse than slightly less complex economies. This provides a theoretical foundation for the finding that countries at high-level of development tend to specialize (e.g., Imbs and Warcziag [82]) and is consistent with the notion that $ECI$ is higher for “small” yet sophisticated and somewhat specialized economies, such as those of Singapore, Switzerland, and Finland [^5] While larger and more diverse economies, like those of Spain and Italy, are not necessarily as complex.. Still, the model predicts a positive correlation between capability endowments and diversity, but through a non-monotonic function, explaining why measures of diversity or concentration are non-ideal estimates of the complexity of an economy.
@@
Third, these results also provide a mean to interpret the structure of the networks of related activities, such as the product space [1], industry space [3], or research space [6]. These networks have been used extensively to model path dependencies and generate measures of export or employment potential [83, 9, 6, 3, 5, 4, 84, 66, 85, 86, 87, 88, 89, 90, 91]. Yet, the structure of these networks differs depending on the data used to generate them. For instance, networks derived from co-export data, are known to have a core composed of densely interconnected activities that are high in complexity surrounded by a periphery of low complexity activities [1]. Research spaces, connecting academic fields based on citations [68] or co-authorships [6], follow a ring structure, with fields connected with a few neighbors and without a clear center [^6] In simple, the ring is: medicine, biology, chemistry, physics, computer science and math, economics, cognitive science, neuroscience, and back to medicine.. While these differences in structure are self-evident, we hitherto lacked a method to explain them based on a mechanistic model. Here, we show how to generate network structures that resemble those observed in the empirical literature by changing the shape of the capability endowment matrices.
@@
Together, these findings help solve some long-standing puzzles in the economic complexity and international development literature by providing a theoretical foundation for economic complexity.
@@ 1.1 Empirical and Theoretical Work in Economic Complexity
Empirical work in economic complexity usually starts with matrices summarizing the geography of many economic activities (e.g., exports by country and product, payroll by city and industry, patents by city and technology, etc.). These rectangular matrices (or bipartite networks) are then used for two things. The first one is to estimate networks of similar activities [8, 9, 6, 1, 83, 66, 3, 92, 93, 91, 7] which are used to estimate the diversification potential of an economy. These measures of “relatedness” have been used to establish the principle that economies are more likely to enter (and less likely to exit) activities that share capabilities with each other [^7] What is know in the specialized literature as The “Principle of Relatedness” [83]..
@@
The second one is to create measures of the value of the portfolio of activities an economy specializes in, known as measures of economic complexity [10, 66, 19, 94, 2, 20, 21, 95]. These measures were also motivated as agnostic estimates of the capabilities available in an economy [10] and are often based on the assumption that high-complexity economies specialize in high-complexity activities. In fact, the economic complexity index or $ECI$, defines the complexity of an economy as the average complexity of the activities it specializes in, and the complexity of an activity as the average complexity of the economies specialized in that activity. [^8] A similar definition was proposed over a decade later by [21]. In their words: “If a country is known to be more capable than another, say the United States (US) versus Bangladesh (BG), then one can identify any good $k$ as more complex than another reference good $k_{0}$ if, relative to the reference good, it is more likely to be exported by the United States than Bangladesh. $\dots$] Conversely, if a good is known to be more complex than another, say medicines (ME) versus men’s underwear (UW), then one can identify any country $i_{1}$ as more capable than another reference country $i_{0}$ if, relative to the reference country, it is more likely to export medicines than underwear. ”. These measures of complexity, in particular $ECI$, enjoy wide adoption in international and regional development circles, as they have been shown to be robust estimators of future economic growth [10, 2, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and of international variations in inequality [29, 30, 31, 32, 33, 37, 44], and emissions [38, 39, 40, 41, 42, 43, 96, 45, 46].
These two strands of literature support the notion that an economy’s pattern of specialization matters for subsequent economic development [97, 98], which has been a key intuition motivating these efforts [^9] This policy intuition is connected to an old debate in development economics, going back to at least Alexander Hamilton’s Report on Manufactures [99], which advocated for the industrialization of the United States, and has been central to the works of scholars such as Rosenstein-Rodan [100, 101], Rostow[102], Hirschman[103], Prebisch[104], Gerschenkron[105], and Balassa[106]. For a discussion on how these different development theories related to economic complexity see [107].. Yet, despite copious empirical work, we still lack an understanding of why the eigenvectors used as measures of economic complexity are good predictors of an economy’s subsequent growth and development.
@@
Theoretical work on economic complexity has focused instead on the construction of models of development and innovation that follow a combinatorial tradition  [108, 109, 10, 110, 56, 111, 112, 113, 114, 115, 116, 117]. This tradition builds on the notion that economies are endowed with capabilities [118, 119, 120, 121, 2, 122], or factors, which activities may or may not require. Since these capabilities are complementary, producing an activity requires the simultaneous presence of many of them. That’s why these theories have been dubbed as the “Lego” or “Scrabble” theory of development. In these models, the ability of economies to produce a product depends on having the right combination of capabilities, like in a proverbial game of scrabble where products are “words” and economies are endowed with “letters.”
@@
Here we build on the combinatorial model introduced by [10], which is a generalization of Kremer’s O-Ring model of development [110], or more precisely the Kremer-Shockley model of productivity, since the same multiplicative formula was introduced in 1957 by William Shockley in a paper explaining differences in productivity among researchers [123].
@@
The Kremer-Shockley model assumes a multi-step production process where the output of an economy is the product of the probabilities that it succeeds at each step. In other words, producing an item in this model requires a sequence of tasks, each of which has a probability of failing. This implies that the output of an economy decays exponentially with the length of the production chain at a rate determined by the probability of succeeding at a task. The key outcome is that economies with higher probabilities of completing a task should specialize in activities requiring multiple steps [^10] See also [124] for an extension of the O-Ring model to trade.. Here we focus on a generalized version of this model, where economies are endowed with probabilities of having a capability (similar to the probability of succeeding at a task in the Kremer-Shockley model) and where activities also differ in the probability of requiring a capability. This allows us to model matrices involving an arbitrary number of economies, activities, and capabilities, while also making the capabilities specific to activities and economies. The resulting matrices, which can be made as large as the ones used in the empirical literature, can be used to create theoretical estimates of the economic complexity eigenvector $ECI$ that we can interpret in connection with the key parameter of the model: the unobservable matrix of capability endowments.
@@
We find that, for a wide variety of model specifications, $ECI$ recovers the probability that an economy is endowed with multiple capabilities, even when these are highly idiosyncratic. We then generalize this finding to a Cobb-Douglas type factor intensity production function and find that the ability of $ECI$ to separate among better and worse endowed economies can be generalized to any shifted production function of the form $Y_{cp}=B+f_{c}g_{p}$ where $f_{c}$ is a general function characterizing an economy $c$ and $g_{p}$ is a general function characterizing an activity $p$.
@@
As in Kremer-Shockley [110, 123], we start from a supply-side model that assumes prices are exogenous and do not provide an explicit model of wages or demand. So, we then embed the single capability model in a short-run equilibrium framework to estimate functions for the implied wages, consumption, and prices. We show that wages increase with the capability endowment of an economy, consumption grows with income, and prices are higher for more complex products (products that are more demanding of the capability). This also allows us to show that our main result–that the economic complexity eigenvector separates high- from low-capability economies–holds after introducing these additional assumptions.
@@
Finally, we use this model to explain the structures of networks of related activities, such as the product space and research space, and show that it is possible to generate networks with a similar structure than the ones observed in the empirical literature by manipulation the capability endowment and capability requirement matrices.
@@
The remainder of the paper is organized as follows. The next section (Section 2) introduces the single-capability model and presents an analytical derivation of the economic complexity index ($ECI$). Section 3 generalizes these results numerically to several versions of a multi-capability model. Section 4 explores additional production functions and Section 6 embeds the model in a short-run equilibrium framework. Section 7 uses the model to explain the structure of networks of related activities, and Section 8 concludes.